OpenStudy (1645323):

Quadratic form of this equation:y = ½(x - 4)2 - 3

OpenStudy (1645323):

@terenzreignz

terenzreignz (terenzreignz):

This isn't a question yet, you know :D

OpenStudy (1645323):

I know I just don't know how to turn the vertex form back into the standard form

OpenStudy (1645323):

Its the vertex form in the question

terenzreignz (terenzreignz):

Oh, well, simple enough ^^ Just expand the square of the binomial

OpenStudy (1645323):

so .5(x-2)(x-2)-3?

terenzreignz (terenzreignz):

Why x-2?

OpenStudy (1645323):

because it adds to -4

terenzreignz (terenzreignz):

No, stay with the actual definition of squaring :D x^2 = x x (x-4)^2 = (x-4)(x-4) Can you expand further?

OpenStudy (1645323):

i don't know at all

terenzreignz (terenzreignz):

FOIL method?

OpenStudy (1645323):

oh so x^2-8x+16

terenzreignz (terenzreignz):

Yes... exactly XD Now, replace (x-4)^2 with that... and then just distribute and simplify.

OpenStudy (1645323):

ok so x^2+-4x+13?

OpenStudy (1645323):

no x^2-4x+5

terenzreignz (terenzreignz):

Almost. You forgot to multiply x^2 with ½ :P

OpenStudy (1645323):

what would it be then .5x^2?

terenzreignz (terenzreignz):

Yup. ^_^

terenzreignz (terenzreignz):

I prefer ½x^2 but whatever floats your boat XD

OpenStudy (1645323):

ok

terenzreignz (terenzreignz):

Now, that wasn't so hard, was it?